Magnesium hydroxide was reacted with hydrochloric acid. The weight of magnesium hydroxide was 1500 grams. What was the weight of magnesium chloride being produced from the reaction?
1
Expert's answer
2020-12-03T14:03:01-0500

Mg (OH) 2 + 2HCl = MgCl2 + 2H2O

n (Mg (OH) 2) = m / M = 1500/58 = 25,86 моль

M (Mg (OH) 2) = 24 + 16 * 2 + 1 * 2 = 58 г / моль

n (MgCl2) = 25,86 моль

т = п * М

M (MgCl2) = 24 * 1 + 35,5 * 2 = 95 г / моль

m (MgCl2) = 25,86 * 95 = 2457 г.

Ответ: 2457 г